OpenStudy (anonymous):

7th grade math please help!

OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (mayankdevnani):

use DISTRIBUTIVE PROPERTY

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that doesnt help much lol

OpenStudy (mayankdevnani):

hope you understand !

OpenStudy (mayankdevnani):

understood ?

OpenStudy (justin528p):

i got u give me a sec to solve it

OpenStudy (mayankdevnani):

WELCOME TO OPEN STUDY ! @justin528p

OpenStudy (mayankdevnani):

@zoejay1

OpenStudy (justin528p):

Simplifying 5(8 + 7) = 5x + 5 * 7 Combine like terms: 8 + 7 = 15 5(15) = 5x + 5 * 7 Multiply 5 * 15 75 = 5x + 5 * 7 Multiply 5 * 7 75 = 5x + 35 Reorder the terms: 75 = 35 + 5x Solving 75 = 35 + 5x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-5x' to each side of the equation. 75 + -5x = 35 + 5x + -5x Combine like terms: 5x + -5x = 0 75 + -5x = 35 + 0 75 + -5x = 35 Add '-75' to each side of the equation. 75 + -75 + -5x = 35 + -75 Combine like terms: 75 + -75 = 0 0 + -5x = 35 + -75 -5x = 35 + -75 Combine like terms: 35 + -75 = -40 -5x = -40 Divide each side by '-5'. x = 8 Simplifying x = 8

OpenStudy (thebossofme191):

basically add 8 and 7 then multiply that by 5 then to get x multiply 5and7 then subtract that from the answer u got by multiplying 5by 15 or 8+7

OpenStudy (mayankdevnani):

don't give direct answers @justin528p you violate CODE OF CONDUCT !

OpenStudy (anonymous):

but he explained how to do it.

OpenStudy (justin528p):

oh o-o sowwy i is new

OpenStudy (mayankdevnani):

@zoejay1 the method i told you is best and short method !

OpenStudy (justin528p):

and she has eh point :)

OpenStudy (mayankdevnani):

@justin528p its a warning ! never do it again

OpenStudy (justin528p):

kk do all exept answer

OpenStudy (justin528p):

:P

OpenStudy (mayankdevnani):

@zoejay1 you didn't reply me !

OpenStudy (justin528p):

hehe still easy solution wit meh:P

OpenStudy (anonymous):

sorry!

OpenStudy (mayankdevnani):

don't worry !

OpenStudy (justin528p):

she didnt have to i showed her how to do it o-o

OpenStudy (mayankdevnani):

I'll tell you short and best method !

OpenStudy (justin528p):

o-o.....

OpenStudy (justin528p):

woah someone wants that medal bad o-o

OpenStudy (mayankdevnani):